A solution


of H2SO4 with a molal concentration of 8.010 m has a density of 1.354 g/mL. What is the


molar concentration of this solution?



1
Expert's answer
2021-02-15T03:25:51-0500

To solve for the molarity, we follow the formula below:

C = n / V

We therefore need to solve for the moles of solute and liters of solution. We can do two steps for this problem:

1. Get the moles of solute from the molality.

2. Determine the volume of the solution using the given density.

For the first step, we can assume that we have 1 kg of the solution to simplify our calculation on the moles of solute. This gives us nH2SO4 = 8.010 (mol)

For the second step, we still use the same assumption for the calculation to be valid:

V = 1 / 1.354 = 0.73855 (L)

Using the values from steps 1 and 2, we have:

C = 8.010 / 0.73855 = 10.85 (M)

Answer: CH2SO4 = 10.85 M.